A thermometer which was not accurately calibrated indicates -0.5°c at the lower fixed point, and the 106°c at the upper fixed point. What temperature does the thermometer register when the temperature is 60°c.


1
Expert's answer
2022-01-24T11:17:09-0500

Let the temperature the thermometer registers be x°

"\frac{60-0.5}{106-0.5}=\frac{x-0}{100-0}"


"x=\frac{59.5×100}{105.5}=56.4°C" ( to 1 d.p.)